# Nadaraya–Watson kernel smoothing as a random energy model

### v1

@@ Abstract
We investigate the behavior of the Nadaraya–Watson kernel smoothing estimator in high dimensions using its relationship to the random energy model and to dense associative memories.
@@ I Introduction
In this note, we explore an approach to characterizing the NW estimator based on the observation that it can be interpreted in terms of a Random Energy Model (REM) [9, 10, 11]  [^1] See Mézard and Montanari [9] for a pedagogical introduction to the REM.. Concretely, we can view (1) as the average of the observable $f(x_{\mu})$ with respect to a distribution over $\mu$ with Boltzmann weights $k(x,x_{\mu})$ that have quenched randomness due to the randomness of the datapoints $x_{\mu}$. For suitable choices of the data distribution, kernel, and target function, it should thus be possible to apply standard approaches to the REM to study the high-dimensional behavior of the NW estimator. The slightly peculiar feature of the problem is that both the energies and the observable depend on the quenched disorder, i.e., on the random data. From the REM perspective, the relevant regime is when $n=e^{\alpha d}$ as $d\to\infty$ for some fixed $\alpha>0$, which would match the sample complexity required by the curse of dimensionality. Moreover, one expects that under suitable conditions the REM associated to the estimator will display a condensation transition: there will be a phase in which all datapoints contribute relatively equally to the estimate, and another in which the distribution is sharply condensed, and only a few points contribute.
@@ II Large deviations analysis for spherical data
We now leverage the REM analogy to analyze the asymptotic behavior of $\hat{f}_{\mathcal{D}}(x)$. For simplicity, we fix a concrete model for the data, kernel, and target function, though many parts of this analysis could be generalized. We assume that the inputs lie on the sphere of radius $\sqrt{d}$, with a fixed test point $x\in\mathbb{S}^{d-1}(\sqrt{d})$ and uniformly-drawn training examples $x_{\mu}\sim\mathcal{U}[\mathbb{S}^{d-1}(\sqrt{d})]$, and take the kernel to be a radial basis function on the sphere, i.e.,
@@
for some inverse bandwidth $\beta>0$. This makes the mapping to the REM quite direct, and is particularly closely related to Lucibello and Mézard [10]’s recent work on dense associative memories (DAMs) [12], where it is shown that the memorization capacity for spherical patterns is determined by the free energy of an auxiliary REM. With these choices, we focus on the simplest class of functions on the sphere with underlying low-dimensional structure: single index models. Such functions take the form
@@
for a vector $w\in\mathbb{S}^{d-1}(\sqrt{d})$ and a scalar link function $g:\mathbb{R}\to\mathbb{R}$. We will often consider positive-homogeneous link functions satisfying $g(\gamma t)=\gamma^{k}g(t)$ for all $\gamma>0$ and some degree $k\geq 0$; among these is of course the linear rectifier $g(t)=\max\{0,t\}$.
We begin by considering the predictions on some fixed test point $x\in\mathbb{S}^{d-1}(\sqrt{d})$. We will argue that there is an explicitly computable $r_{\ast}=r_{\ast}(\alpha,\beta)$, bounded from below by zero and from above by 1, such that we have the pointwise asymptotic
@@
This potential will consist of an energetic contribution $\beta t$ from the kernel, plus an entropic contribution resulting from re-writing the sum (6) in terms of the density of states with a given $(t,q)$. We will follow the standard random energy model analysis to obtain an asymptotic of this form, closely following the related work of Lucibello and Mézard [10]. Pedagogical accounts of this analysis can also be found in the textbook of Mézard and Montanari [9], or in the original paper of Derrida [11].
@@ III Mean-squared generalization error
The asymptotic (28) gives us a prediction for the absolute deviation $|f(x)-\hat{f}_{\mathcal{D}}(x)|$ for a (typical) test point, but one would also want to compute the asymptotics of the mean-squared generalization error. For the single-index models we consider, this reduces to a scalar average over the distribution of $\rho$ induced by the randomness in $x$. Asymptotically, $\rho$ should be approximately Gaussian with variance $1/d$. Given that this distribution is manifestly dimension-dependent, we are confronted with the question of the scale at which we should measure the generalization error, and the question of whether the finite-size corrections to the simple asymptotic for a fixed $x$ must be taken into account in order to accurately predict the mean-squared generalization error.
@@ V Numerical experiments
VI Conclusions
@@
There are many problems which we leave open for future inquiry. First, though we have contented ourselves with simple asymptotics and comparisons to limited numerics, this analysis could in principle be made entirely quantitative in the sense of explicit error bounds as a function of $d$. Some large deviations results for the NW estimator in fixed dimension $d$ as $n\to\infty$ are known [13, 3], but to our knowledge rigorous results when $d$ and $n$ tend to infinity together are lacking. Given the close relationship of the settings, rigorous bounds on the finite-size effects for the setting considered here would immediately imply similar bounds for memorization of spherical patterns in DAMs [10]. Second, we have specialized to the setting of spherical data, a radial basis function kernel, and a single-index target; in principle all of these assumptions could be relaxed. Extending our results to anisotropic data would be of particular interest, as introducing a preferred direction at some specified angle with the latent vector $w$ would likely alter the condensation threshold due to the local enhancement of data density. A natural starting point is thus to consider data drawn from a von Mises-Fisher distribution, which would require one to track three overlaps between the latent vector $w$, the test point $x$, and the mean direction $\eta$.
Finally, we comment on the broader context of our work. In the last few years, substantial effort has been devoted to seeking precise characterizations of how various learning algorithms behave in high dimensions. As mentioned in the introduction to this note, a key achievement of this program has been a sharp understanding of kernel ridge regression [4, 5, 6, 2, 7]. This essay represents a small step towards a similarly detailed understanding of qualitatively different regression algorithms. Thus, in closing, we echo Jamie Simon’s call to arms [14]: let’s solve more learning rules!
@@ Acknowledgements
We thank Jamie Simon for posing the question that inspired this note [8], and thank Alexander Atanasov, Blake Bordelon, Benjamin Ruben, and especially Sabarish Sainathan for useful discussions. JAZV further thanks Dmitry Krotov for useful discussions regarding [10]. JAZV is supported by a Junior Fellowship from the Harvard Society of Fellows. JAZV and CP were supported by NSF Award DMS-2134157 and NSF CAREER Award IIS-2239780. CP is further supported by a Sloan Research Fellowship. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. The computations in this paper were run on the FASRC Cannon cluster supported by the FAS Division of Science Research Computing Group at Harvard University.

### v2

@@ Abstract
Precise asymptotics have revealed many surprises in high-dimensional regression. These advances, however, have not extended to perhaps the simplest estimator: direct Nadaraya-Watson (NW) kernel smoothing. Here, we describe how one can use ideas from the analysis of the random energy model (REM) in statistical physics to compute sharp asymptotics for the NW estimator when the sample size is exponential in the dimension. As a simple starting point for investigation, we focus on the case in which one aims to estimate a single-index target function using a radial basis function kernel on the sphere. Our main result is a pointwise asymptotic for the NW predictor, showing that it re-scales the argument of the true link function. Our work provides a first step towards a detailed understanding of kernel smoothing in high dimensions.
@@ I Introduction
In this note, we explore an approach to characterizing the NW estimator based on the observation that it can be interpreted in terms of a Random Energy Model (REM) [9, 10, 11]. [^1] See Mézard and Montanari [9] for a pedagogical introduction to the REM. Concretely, we can view (1) as the average of the observable $f(x_{\mu})$ with respect to a distribution over $\mu$ with Boltzmann weights $k(x,x_{\mu})$ that have quenched randomness due to the randomness of the datapoints $x_{\mu}$. For suitable choices of the data distribution, kernel, and target function, it should thus be possible to apply standard approaches to the REM to study the high-dimensional behavior of the NW estimator. The slightly peculiar feature of the problem is that both the energies and the observable depend on the quenched disorder, i.e., on the random data. From the REM perspective, the relevant regime is when $n=e^{\alpha d}$ as $d\to\infty$ for some fixed $\alpha>0$, which would match the sample complexity required by the curse of dimensionality. Moreover, one expects that under suitable conditions the REM associated to the estimator will display a condensation transition: there will be a phase in which all datapoints contribute relatively equally to the estimate, and another in which the distribution is sharply condensed, and only a few points contribute.
@@ II Large deviations analysis for spherical data
We now leverage the REM analogy to analyze the asymptotic behavior of $\hat{f}_{\mathcal{D}}(x)$. For simplicity, we fix a concrete model for the data, kernel, and target function, though many parts of this analysis can be generalized. We assume that the inputs lie on the sphere of radius $\sqrt{d}$, with a fixed test point $x\in\mathbb{S}^{d-1}(\sqrt{d})$ and uniformly-drawn training examples $x_{\mu}\sim\mathcal{U}[\mathbb{S}^{d-1}(\sqrt{d})]$, and take the kernel to be a radial basis function on the sphere, i.e.,
@@
for some inverse bandwidth $\beta>0$. This makes the mapping to the REM quite direct, and is particularly closely related to Lucibello and Mézard [10]’s recent work on dense associative memories (DAMs) [12], where it is shown that the memorization capacity for spherical patterns is determined by the free energy of an auxiliary REM.
With these choices, we focus on the simplest class of functions on the sphere with underlying low-dimensional structure: single index models. Such functions take the form
@@
for a vector $w\in\mathbb{S}^{d-1}(\sqrt{d})$ and a scalar link function $g:\mathbb{R}\to\mathbb{R}$. For now, we will assume that the function $g$ is independent of the dimension $d$; we relax that assumption in Section VI. We will often consider positive-homogeneous link functions satisfying $g(\gamma t)=\gamma^{k}g(t)$ for all $\gamma>0$ and some degree $k\geq 0$; among these is of course the linear rectifier $g(t)=\max\{0,t\}$.
We begin by considering the predictions on some fixed test point $x\in\mathbb{S}^{d-1}(\sqrt{d})$. For dimension-independent link functions, we will argue that there is an explicitly computable $r_{\ast}=r_{\ast}(\alpha,\beta)$, bounded from below by zero and from above by 1, such that we have the pointwise asymptotic
@@
This potential will consist of an energetic contribution $\beta t$ from the kernel, plus an entropic contribution resulting from re-writing the sum (6) in terms of the density of states with a given $(t,q)$. We will follow the standard random energy model analysis to obtain an asymptotic of this form, closely following the related work of Lucibello and Mézard [10]. [^2] As we will invoke results from [10] without proof, we provide a brief dictionary of notation in Appendix A. Pedagogical accounts of this analysis can also be found in the textbook of Mézard and Montanari [9], or in the original paper of Derrida [11].
@@ III Mean-squared generalization error
The asymptotic (28) gives us a prediction for the absolute deviation $|f(x)-\hat{f}_{\mathcal{D}}(x)|$ for a (typical) test point, but one would also want to compute the asymptotics of the mean-squared generalization error. For the single-index models we consider, this reduces to a scalar average over the distribution of $\rho$ induced by the randomness in $x$. Asymptotically, $\rho$ should be approximately Gaussian with variance $1/d$. Given that this distribution is manifestly dimension-dependent, we are confronted with the question of the scale at which we should measure the generalization error, and the question of whether the finite-size corrections to the simple asymptotic for a fixed $x$ must be taken into account in order to accurately predict the mean-squared generalization error. An important effect to consider in this case is the fact that the distribution of fluctuations of the weights in the REM depends strongly on the phase: in the condensed phase, fluctuations become heavy-tailed [11, 13].
@@ V Numerical experiments
VI Large target functions
Up to this point, we have assumed that the target link function $g$ does not scale with dimension, and thus is small in the sense that it does not affect the maximization of the potential in (18). If the target function is large enough to change the location of the maximum, the relationship between the true function and the NW estimator is more complicated asymptotically than a simple renormalization of the argument. This can be seen by considering the case in which the target function is itself a radial basis function on the sphere,
$g(\langle w,x\rangle/d)=e^{\gamma\langle w,x\rangle},$ (38)
where $\gamma>0$ is the inverse bandwidth of the target. Then, to leading exponential order, we are interested in the difference of the free energies of two REMs at different inverse temperatures. Concretely, we consider
$\frac{1}{d}\log\hat{f}_{\mathcal{D}}(x)=\frac{1}{d}\log\sum_{\mu=1}^{n}e^{\langle\beta x+\gamma w,x_{\mu}\rangle}-\frac{1}{d}\log\sum_{\mu=1}^{n}e^{\beta\langle x,x_{\mu}\rangle}.$ (39)
From this point, we could either adapt the pointwise large deviations analysis of Section II, or leverage the fact that this is now actually identical to the standard REM computation done in Lucibello and Mézard [10]. By either route, one finds that
$\frac{1}{d}\log\hat{f}_{\mathcal{D}}(x)\sim\phi_{\ast}(\sqrt{\beta^{2}+\gamma^{2}+2\beta\gamma\rho})-\phi_{\ast}(\beta),$ (40)
where by a minor abuse of notation we write $\phi_{\ast}(\beta)=\phi(r_{\ast}(\beta))$ for the value of the radial potential
$\phi(r_{\ast})=\alpha+\beta r_{\ast}+\frac{1}{2}\log(1-r_{\ast}^{2})$ (41)
maximized with respect to $r\in(0,\sqrt{1-e^{-2\alpha}}]$ for a particular value $\beta$ of the bandwidth.
This result shows that if the target function is allowed to grow with dimension then the pointwise asymptotic behavior of the NW estimator is not always simple multiplicative renormalization. In particular, this follows from the fact that $\rho$ enters the asymptotic only through $\sqrt{\beta^{2}+\gamma^{2}+2\beta\gamma\rho}$, which unless $\gamma$ is perturbatively small is far from linear in $\rho$. Moreover, in this case the choice of test point can affect whether the estimator is in the condensed phase, which contrasts with what we found in the case of $n$-independent link functions. In particular, if both $\sqrt{\beta^{2}+\gamma^{2}+2\beta\gamma\rho}$ and $\beta$ are above the condensation threshold $e^{2\alpha}\sqrt{1-e^{-2\alpha}}$, then this asymptotic predicts that $\frac{1}{d}\log\hat{f}_{\mathcal{D}}(x)$ will vanish to leading order.
VII Conclusions
@@
There are many problems which we leave open for future inquiry. Though we have contented ourselves with simple asymptotics and comparisons to limited numerics, this analysis could in principle be made entirely quantitative in the sense of explicit error bounds as a function of $d$. Some large deviations results for the NW estimator in fixed dimension $d$ as $n\to\infty$ are known [14, 3], but to our knowledge rigorous results when $d$ and $n$ tend to infinity together are lacking. Given the close relationship of the settings, rigorous bounds on the finite-size effects for the setting considered here would immediately imply similar bounds for memorization of spherical patterns in DAMs [10].
We have specialized to the setting of spherical data, radial basis function kernels, and a single-index target; in principle all of these assumptions could be relaxed. Extending our results to anisotropic data would be of particular interest, as introducing a preferred direction at some specified angle with the latent vector $w$ would likely alter the condensation threshold due to the local enhancement of data density. A natural starting point is thus to consider data drawn from a von Mises-Fisher distribution, which would require one to track three overlaps between the latent vector $w$, the test point $x$, and the mean direction $\eta$. Extensions to multi-index targets would be of a similar flavor: so long as the number of relevant directions remains finite, one would have to characterize the joint large deviations of the overlap between the input and each direction. This would make the computation more complicated, but at least at the level of a pointwise asymptotic would not introduce new technical challenges. Finally, our approach could be adapted to other kernels provided that one accounts for how their variances scale with dimension. This issue is closely related to classical studies of REM-type models with general energy distributions, as analyzed in detail by Ben Arous et al. [13].
Finally, we comment on the broader context of our work. In the last few years, substantial effort has been devoted to seeking precise characterizations of how various learning algorithms behave in high dimensions. As mentioned in the introduction to this note, a key achievement of this program has been a sharp understanding of kernel ridge regression [4, 5, 6, 2, 7]. It is interesting to contrast the multiplicative renormalization of the argument of the link function observed here to the case of KRR, for which the randomness in the data renormalizes the ridge parameter itself [6]. This essay represents a small step towards a similarly detailed understanding of qualitatively different regression algorithms. In concurrent work posted to the arXiv shortly after the initial appearance of this note, Biroli and Mézard [15] have performed a detailed analysis of kernel density estimation that similarly leverages an analogy to the REM. The time is therefore ripe for a full asymptotic description of kernel smoothing methods. Thus, in closing, we echo Jamie Simon’s call to arms [16]: let’s solve more learning rules!
@@ Acknowledgements
We thank Jamie Simon for posing the question that inspired this note [8], and thank Alexander Atanasov, Blake Bordelon, Benjamin Ruben, and especially Sabarish Sainathan for useful discussions. JAZV further thanks Dmitry Krotov for useful discussions regarding [10].
Funding acknowledgements
JAZV is supported by a Junior Fellowship from the Harvard Society of Fellows. JAZV and CP were supported by NSF Award DMS-2134157 and NSF CAREER Award IIS-2239780. CP is further supported by a Sloan Research Fellowship. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. The computations in this paper were run on the FASRC Cannon cluster supported by the FAS Division of Science Research Computing Group at Harvard University.
Author contributions
JAZ-V conceived the project, performed all research, and wrote the paper. CP supervised the project and contributed to review and editing.
Appendix A Table of notation
Here, we provide a brief table of notation, with translations to the notation of Lucibello and Mézard [10] for the reader’s convenience. We direct the reader in particular to Appendix C of Lucibello and Mézard [10], which contains their analysis of the rate function for spherical data, referred to as “patterns” in the context of associative memory models.
Quantity Our notation Lucibello and Mézard [10]
dimension $d$ $N$
number of training examples (“patterns”) $n=e^{\alpha d}$ $P=e^{\alpha N}$
training examples (“patterns”) $\mathbf{x}_{\mu}$ $\bm{\xi}^{\mu}$
rate function $\zeta$ $\zeta_{\rho}$
REM free energy density $\beta^{-1}\phi$ $\phi$
inverse temperature/inverse bandwidth $\beta$ $\lambda$